The functions 2 = -e 3t 2 Y = 3e -3t solve the system Dx +(D + 2)y=0. (D - 3x – 2y = 0 True False
Answer #1

Given system is ܝܘ The functions .3t DX + (D+2) = 0 5 Da (D-3) x-2y = 0 (2) -37 will be a solun of the system of they satisfi